# Convergence of Unadjusted Langevin in High Dimensions: Delocalization of Bias

### v1

@@ 1.1. Motivation
[…]
@@
In this paper, we argue that the mismatch between theory and practice is due, in large part, to the metric used to measure convergence. Standard metrics such as $W_{2}$ measure accuracy of the entire distribution, but practitioners are often interested in averages of functions involving a relatively small set of variables. Based on this observation, we aim to investigate the convergence behavior of the algorithm under alternative metrics designed to characterize the accuracy of low-dimensional marginals. We identify a new “delocalization” phenomenon for the bias of the unadjusted Langevin algorithm to show that, even in high-dimensional settings, the step size may not need to be very small if the quantities of interest depend on low-dimensional marginals only. Consequently, the iteration complexity required to achieve a bounded error exhibits a benign dependence on dimension $d$.
@@ 1.2. Literature review
References [2] and [14] studied target distributions for which larger stepsizes are possible in high-dimensions if the observable of interest has Lipschitz constant that scales as $1/\sqrt{d}$, such as the averaged observable $f(x)=\frac{1}{d}\sum_{i=1}^{d}\Phi(x^{(i)})$ with $\ell^{2}$-Lipschitz $\Phi$. In these cases, a constant step size is sufficient to achieve a bounded error regardless of dimensionality. Our results go beyond these studies by establishing bounds that are not limited to averaged observables. Our analysis implies that the algorithm could achieve a bounded error for low-dimensional marginals in any dimension using a step size that is dimension-independent up to logarithmic terms.
@@ 1.3.1. A new metric to measure convergence
where $|\cdot|_{\infty}$ is the $\ell^{\infty}$ norm of a vector and $\Pi(\mu,\nu)$ represents the set of measures in the joint space $\mathbb{R}^{d}\times\mathbb{R}^{d}$ that have marginals $\mu,\nu$.
@@ 1.3.2. Positive and negative examples: delocalization effect
Consider $\pi\propto\exp(-V)$ where $V(x)=\sum_{i=1}^{d}V_{i}(x^{(i)})$ satisfies $\alpha\leq\nabla^{2}V_{i}\leq\beta$. Then, for $h\leq\alpha/\beta^{2}$, it holds that
@@
Here $(b)$ is the one-step discretization error which is bounded by $O(\beta h^{3/2}\sqrt{\log(2d)})$ given $h\leq\alpha/\beta^{2}$. And $(a)=\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}-h(\nabla V(X_{kh})-\nabla V(Y_{kh}))|_{\infty}^{2}]}\leq(1-\alpha h)\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}\leq\exp(-\alpha h)\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}$ where we apply the strong convexity of $V_{i}$ to bound each coordinate of the vector yielding the final $\ell^{\infty}$ norm bound. This shows that there is a one-step contraction in the $\ell^{\infty}$ norm. We then couple the distribution of $X_{kh}$ and $Y_{kh}$ such that $\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}=W_{2,\ell^{\infty}}(\rho_{kh},\pi)$. Using the definition of the $W_{2,\ell^{\infty}}$ norm, we get
@@
Consider $\pi=\rho^{\otimes d}$ where $\rho$ is a one-dimensional centered log concave distribution for which the biased distribution $\rho_{h}$ obtained by the corresponding one dimensional unadjusted Langevin has nonzero mean [^1] For example, Proposition 3.1 can be used to show that $\rho$ can be taken to be a univariate mixture of Gaussians $p\mathcal{N}(\mu_{1},1)+(1-p)\mathcal{N}(\mu_{2},1)$ with $p\mu_{1}+(1-p)\mu_{2}=0$, $20<p<1/$ 2, and $|\mu_{1}-\mu_{2}|$ sufficiently small., so that their mean differs by $\delta>0$. Consider the observable $f(x)=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}x^{(i)}$ for $x\in\mathbb{R}^{d}$. It holds that
@@
Example 1.4 shows that for the rotated product measure $\tilde{\pi}$, the $W_{2,\ell^{\infty}}$ bias is of order $\sqrt{d}$. This indicates that the bias is not declocalized, but concentrated on one specific dimension. Thus, the delocalization effect of the bias over marginals is a delicate phenomenon and does not hold universally. We note that Example 1.4 is characterized by strong, dense interactions between coordinates after the rotation.
@@ 1.3.3. Strongly log-concave target distributions with sparse interactions
Some $i$-th variable: $x^{(i)}$1st layer neighborhood: $\mathsf{N}_{1}(i)$2nd layer neighborhood: $\mathsf{N}_{2}(i)$$V(x)=\sum_{i=1}^{d}V_{i}(\mathcal{X}^{i})$, where $\mathcal{X}^{i}=\{x^{(j)}:j\in\mathsf{N}_{1}(i)\}$,and $V_{i}$ is a function depending on variables in $\mathcal{X}^{i}$parsity parameter $s_{k}:=\sup_{1\leq i\leq d}|\mathsf{N}_{k}(i)|$For this example, $s_{k}=O(k^{2})$
@@
Let Assumption 1.1 hold. Suppose $V$ further satisfies the sparsity conditions illustrated in Figure 1 and rigorously formulated in Section 2.1 below with the sparsity parameter $s_{k}\leq C(k+1)^{n}$, then
@@ 1.5. Notations
We write $A=O(B)$ or $A\lesssim B$ to mean that there exists a constant $C$ independent of $\alpha,\beta,h,d$ such that $A\leq CB$. On the other hand, $A=\Omega(B)$ or $A\gtrsim B$ means that there exists a constant $C$ independent of $\alpha,\beta,h,d$ such that $A\geq CB$. We use $|\cdot|_{\infty}$ and $|\cdot|_{2}$ to represent the $\ell^{\infty}$ and $\ell^{2}$ norms for vectors and matrices. When applied to matrices, they stand for the corresponding operator norms. We use $\preceq$ for the Loewner order such that if $M\preceq N$ where $M,N$ are symmetric matrices, then the matrix $N-M$ is positive semi-definite.
@@ 2.1. Sparse graphical models
Consider an undirected graph $G$ with $d$ nodes, labeled by $1\leq i\leq d$. If there is an edge between node $i$ and $j$, we write $i\sim j$, meaning that the two nodes are connected. We denote by $\mathsf{N}(i)$ the neighborhood of the node $i$, which is a set of nodes connected to $i$. Here the neighborhood relationship is symmetric, and without loss of generality, we assume $i\sim i$ for all $1\leq i\leq d$. Furthermore, we define recursively that $\mathsf{N}_{k}(i)=\{1\leq j\leq d:\exists\ \ell\in\mathsf{N}_{k-1}(i),\text{ such that }j\sim\ell\}$, for $k\geq 2$. Here $\mathsf{N}_{1}(i):=\mathsf{N}(i)$. Denote the cardinality of the set $\mathsf{N}_{k}(i)$ by $|\mathsf{N}_{k}(i)|$ and let $\max_{1\leq i\leq d}|\mathsf{N}_{k}(i)|=s_{k}$. We also write $s=s_{1}$.
@@
We can also understand $G$ as a factor graph, where $F=\{\mathcal{X}^{i},1\leq i\leq d\}$ represents the factor vertices. By definition, the potential $V$ described above exhibits sparse interaction when the graph is sparse.
@@ 2.2. Convergence bound for unadjusted Langevin
Let Assumptions 1.1 and 2.1 hold. Assume $0\leq h\leq 1/\beta$ and denote $q=\exp(-h\alpha)$. Let $r_{i}=\lceil e^{2}ih\beta+\log\sqrt{d}\rceil$ for $i\in\mathbb{N}$. Then, for any $N\in\mathbb{N}$ such that $12q^{N}\sqrt{s_{r_{N}}}<$ 1, we have the bound
@@
$:=\frac{2\beta\sum_{i=1}^{N}q^{i-1}\sqrt{s_{r_{i}}}}{1-2q^{N}\sqrt{s_{r_{N}}}}\left(h^{2}\sqrt{\mathbb{E}_{\pi}[|\nabla V|_{\infty}^{2}]}+3h^{3/2}\sqrt{\log(2d)}\right)\,.$
@@
$\sqrt{\mathbb{E}[|X_{(k+1)h}-Y_{(k+1)h}|_{\infty}^{2}]}\leq\sqrt{\mathbb{E}[|H_{k}H_{k-1}(X_{(k-1)h}-{Y}_{(k-1)h})|_{\infty}]}+\mathrm{error}(2)\,,$
@@
$\sqrt{\mathbb{E}[|X_{(k+N)h}-Y_{(k+N)h}|_{\infty}^{2}]}\leq\sqrt{\mathbb{E}[|H_{k+N-1}H_{k+N-2}\cdots H_{k}(X_{kh}-{Y}_{kh})|_{\infty}]}+\mathrm{error}(N)\,.$
@@
$|H_{k+N-1}H_{k+N-2}\cdots H_{k}|_{\infty}\leq\sqrt{d}|H_{k+N-1}H_{k+N-2}\cdots H_{k}|_{2}\leq\sqrt{d}\exp(-N\alpha h)\,.$
@@
for any $k,N\in\mathbb{N}$. When $12q^{N}\sqrt{s_{r_{N}}}<$ 1, we can let $k\to\infty$ and get the final bound in (2.2). The complete proof for Theorem 2.2 can be found in Appendix C. ∎
@@
(2.4) $\sqrt{\mathbb{E}_{\pi}[|\nabla V|_{\infty}^{2}]}\leq 2\frac{\beta}{\sqrt{\alpha}}\sqrt{\log(2d)}\,.$
@@
When $V$ is strongly convex, by the Bakry-Émery criterion, $\pi$ satisfies the log-Sobolev inequality with constant $1/\alpha$. Using this fact and Herbst’s argument leads to the bound (2.4). The complete proof can be found in Appendix C.3. ∎
@@
Let Assumptions 1.1 and 2.1 hold with $s_{k}\leq C(k+1)^{n}$ for any $k\in\mathbb{N}$, where $C>0$ is a generic constant and $n\in\mathbb{N}$. Then for $h\leq\frac{\alpha}{\beta^{2}}$, we have
@@
The current upper bound in Theorem 2.4 involves $\sqrt{h}$. This is in alignment with the $W_{2}$ bound proved in the literature (under Assumption 1.1) based on coupling arguments where such square root dependence manifests. The second part of the bound, $\sqrt{h\log(2d)}O\left(\frac{\beta}{\alpha}\sqrt{d}\right)$, is consistent, up to logarithmic terms in $d$, with the established result for the $W_{2}$ bound. In fact, using the $W_{2}$ bound, we have that $W_{2,\ell^{\infty}}(\pi_{h},\pi)\leq W_{2}(\pi_{h},\pi)=O(\frac{\beta}{\alpha}\sqrt{dh})$. $\lozenge$

### v2

@@ 1.1. Motivation
The unadjusted Langevin algorithm (1.2) is biased, meaning that even when $k\to\infty$, the distribution $\rho_{kh}$ will not converge to the exact target distribution $\pi$ for any finite step size $h$. For strongly log-concave distributions, $\rho_{kh}$ converges in the $W_{2}$ metric to the stationary distribution of (1.2), $\pi_{h}$, at a rate that is independent of dimension (see, e.g., [18]). Consequently, the dependence of the error $W_{2}(\rho_{kh},\pi)$ on dimension, $d$, is completely determined by the bias $W_{2}(\pi_{h},\pi)$. This bias depends on the step size $h$, which, according to state-of-the-art analyses, needs to scale proportionally to $1/d$ or $1/\sqrt{d}$ to achieve a bounded bias for any $d$. Since the iteration complexity typically scales as $1/h$ (up to logarithmic factors), this small step size implies a computational cost of order $\sqrt{d}$ or $d$.
@@
In this paper, we argue that the mismatch between theory and practice is due, in large part, to the metric used to measure convergence. Standard metrics such as $W_{2}$ measure accuracy of the entire distribution, but practitioners are often interested in averages of functions involving a relatively small set of variables. Indeed, both statistical and physical models often require the inclusion of many latent variables (e.g. solvent variables in molecular dynamics simulations) that are themselves of no direct interest.
Based on these observations, we aim to investigate the convergence behavior of the algorithm under alternative metrics designed to characterize the accuracy of low-dimensional marginals. We identify a new “delocalization” phenomenon for the bias of the unadjusted Langevin algorithm to show that, even in high-dimensional settings, the step size may not need to be very small if the quantities of interest depend on low-dimensional marginals only. Consequently, the iteration complexity required to achieve a bounded error exhibits a benign dependence on dimension $d$.
@@ 1.2. Literature review
References [3] and [15] studied target distributions for which larger stepsizes are possible in high-dimensions if the observable of interest has Lipschitz constant that scales as $1/\sqrt{d}$, such as the averaged observable $f(x)=\frac{1}{d}\sum_{i=1}^{d}\Phi(x^{(i)})$ with $\ell^{2}$-Lipschitz $\Phi$. In these cases, a constant step size is sufficient to achieve a bounded error regardless of dimensionality.
@@ 1.3.1. A new metric to measure convergence
where $|\cdot|_{\infty}$ is the $\ell^{\infty}$ norm of a vector and $\Pi(\mu,\nu)$ represents the set of measures in the joint space $\mathbb{R}^{d}\times\mathbb{R}^{d}$ that have marginals $\mu,\nu$. We also use $|M|_{\infty}$ to denote the $\ell^{\infty}\to\ell^{\infty}$ operator norm of a matrix $M\in\mathbb{R}^{d\times d}$ throughout this article:
$|M|_{\infty}=\sup_{x\in\mathbb{R}^{d}:x\neq 0}\frac{|Mx|_{\infty}}{|x|_{\infty}}\,.$
Similarly, $|\cdot|_{2}$ stands for the $\ell^{2}$ norm and, when applied to matrices, to the corresponding operator norm.
@@
We note that Langevin dynamics is rotation invariant and not sensitive to the choice of basis, as is the standard $W_{2}$ metric. However, a distinct feature of the introduced $W_{2,\ell^{\infty}}$ metric is its dependence on the coordinate system. In fact, as we will see later, the delocalization effect studied in this paper is coordinate dependent, which also justifies the use of the coordinate-dependent $W_{2,\ell^{\infty}}$ metric.
@@ 1.3.2. Positive and negative examples: delocalization effect
Consider $\pi\propto\exp(-V)$ where $V(x)=\sum_{i=1}^{d}V_{i}(x^{(i)})$ satisfies $\alpha\leq\nabla^{2}V_{i}\leq\beta$. Then, for $h\leq 1/\beta$, it holds that
@@
Here $(b)$ is the one-step discretization error which is bounded by $O(\beta h^{3/2}\sqrt{\log(2d)})$ given $h\leq 1/\beta$. And $(a)=\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}-h(\nabla V(X_{kh})-\nabla V(Y_{kh}))|_{\infty}^{2}]}\leq(1-\alpha h)\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}\leq\exp(-\alpha h)\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}$ where we apply the strong convexity of $V_{i}$ to bound each coordinate of the vector yielding the final $\ell^{\infty}$ norm bound. This shows that there is a one-step contraction in the $\ell^{\infty}$ norm. We then couple the distribution of $X_{kh}$ and $Y_{kh}$ such that $\sqrt{\mathbb{E}[|X_{kh}-Y_{kh}|_{\infty}^{2}]}=W_{2,\ell^{\infty}}(\rho_{kh},\pi)$. Using the definition of the $W_{2,\ell^{\infty}}$ norm, we get
@@
Consider $\pi=\rho^{\otimes d}$ where $\rho$ is a one-dimensional centered log concave distribution for which the biased distribution $\rho_{h}$ obtained by the corresponding one dimensional unadjusted Langevin has nonzero mean [^1] For example, Proposition 3.1 can be used to show that $\rho$ can be taken to be a univariate mixture of Gaussians $p\mathcal{N}(\mu_{1},1)+(1-p)\mathcal{N}(\mu_{2},1)$ with $p\mu_{1}+(1-p)\mu_{2}=0$, $0<p<1/2$, and $|\mu_{1}-\mu_{2}|$ sufficiently small., so that their mean differs by $\delta>0$. Consider the observable $f(x)=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}x^{(i)}$ for $x\in\mathbb{R}^{d}$. It holds that
@@
Example 1.4 shows that for the rotated product measure $\tilde{\pi}$, the $W_{2,\ell^{\infty}}$ bias is of order $\sqrt{d}$. This indicates that the bias is not delocalized, but concentrated on one specific dimension. Thus, the delocalization effect of the bias over marginals is a delicate phenomenon and does not hold universally. We note that Example 1.4 is characterized by strong, dense interactions between coordinates after the rotation.
We note that in Example 1.4, the delocalization effect holds for the invariant distribution $\pi$ but not for the rotated $\tilde{\pi}$, revealing that the delocalization phenomenon is coordinate dependent. The coordinate-dependent $W_{2,\ell^{\infty}}$ metric enables the characterization of this phenomenon.
@@ 1.3.3. Strongly log-concave target distributions with sparse interactions
Some $i$-th variable: $x^{(i)}$1st layer neighborhood: $\mathsf{N}_{1}(i)$2nd layer neighborhood: $\mathsf{N}_{2}(i)$$V(x)=\sum_{i=1}^{d}V_{i}(\mathcal{X}^{i})$, where $\mathcal{X}^{i}=\{x^{(j)}:j\in\mathsf{N}_{1}(i)\}$,and $V_{i}$ is a function depending on variables in $\mathcal{X}^{i}$Sparsity parameter $s_{k}:=\sup_{1\leq i\leq d}|\mathsf{N}_{2k}(i)|$For this example, $s_{k}=O(k^{2})$
@@
Let Assumption 1.1 hold. Suppose $V$ further satisfies the sparsity conditions illustrated in Figure 1 and rigorously formulated in Section 2.1 below with the sparsity parameter $s_{k}\leq C(k+1)^{n}$, then for $h\leq 1/\beta$,
@@ 1.5. Notations
We write $A=O(B)$ or $A\lesssim B$ to mean that there exists a constant $C$ independent of $\alpha,\beta,h,d$ such that $A\leq CB$. On the other hand, $A=\Omega(B)$ or $A\gtrsim B$ means that there exists a constant $C$ independent of $\alpha,\beta,h,d$ such that $A\geq CB$. We use $|\cdot|_{\infty}$ and $|\cdot|_{2}$ to represent the $\ell^{\infty}$ and $\ell^{2}$ norms for vectors and matrices. When applied to matrices, they stand for the corresponding operator norms, as in Section 1.3.1. We use $\preceq$ for the Loewner order such that if $M\preceq N$ where $M,N$ are symmetric matrices, then the matrix NMN-M is positive semi-definite.
@@ 2.1. Sparse graphical models
Consider an undirected graph $G$ with $d$ nodes, labeled by $1\leq i\leq d$. If there is an edge between node $i$ and $j$, we write $i\sim j$, meaning that the two nodes are connected. We denote by $\mathsf{N}(i)$ the neighborhood of the node $i$, which is a set of nodes connected to $i$. Here the neighborhood relationship is symmetric, and without loss of generality, we assume $i\sim i$ for all $1\leq i\leq d$. Furthermore, we define recursively that $\mathsf{N}_{k}(i)=\{1\leq j\leq d:\exists\ \ell\in\mathsf{N}_{k-1}(i),\text{ such that }j\sim\ell\}$, for $k\geq 2$. Here $\mathsf{N}_{1}(i):=\mathsf{N}(i)$. Denote the cardinality of the set $\mathsf{N}_{k}(i)$ by $|\mathsf{N}_{k}(i)|$ and let $\max_{1\leq i\leq d}|\mathsf{N}_{2k}(i)|=s_{k}$. We also write $s=s_{1}$.
@@
We can also understand $G$ as a factor graph, where $F=\{\mathcal{X}^{i},1\leq i\leq d\}$ represents the factor vertices. By definition, the potential $V$ described above exhibits sparse interaction when the graph is sparse. We observe that, under Assumption 2.1, the number of non-zero entries in each column of $\nabla^{2}V(x)$ is bounded by $s_{1}$. More generally, when multiplying $k$ such Hessians, the number of non-zero entries per column is bounded by $s_{k}$, a fact that plays a key role in our subsequent analysis.
@@ 2.2. Convergence bound for unadjusted Langevin
Let Assumptions 1.1 and 2.1 hold. Assume $0\leq h\leq 1/\beta$ and denote $q=\exp(-h\alpha)$. Let $r_{i}=\lceil e^{2}ih\beta+\log\sqrt{d}\rceil$ for $i\in\mathbb{N}$. Then, for any $N\in\mathbb{N}$ such that $2q^{N}\sqrt{s_{r_{N}}}<1$, we have the bound
@@
$=\frac{2\beta\sum_{i=1}^{N}q^{i-1}\sqrt{s_{r_{i}}}}{1-2q^{N}\sqrt{s_{r_{N}}}}\left(h^{2}\sqrt{\mathbb{E}_{\pi}[|\nabla V|_{\infty}^{2}]}+3h^{3/2}\sqrt{\log(2d)}\right)\,.$
@@
$\sqrt{\mathbb{E}[|X_{(k+1)h}-Y_{(k+1)h}|_{\infty}^{2}]}\leq\sqrt{\mathbb{E}[|H_{k}H_{k-1}(X_{(k-1)h}-{Y}_{(k-1)h})|_{\infty}^{2}]}+\mathrm{error}(2)\,$,
@@
$\sqrt{\mathbb{E}[|X_{(k+N)h}-Y_{(k+N)h}|_{\infty}^{2}]}\leq\sqrt{\mathbb{E}[|H_{k+N-1}H_{k+N-2}\cdots H_{k}(X_{kh}-{Y}_{kh})|_{\infty}^{2}]}+\mathrm{error}(N)\,.$
@@
$|H_{k+N-1}H_{k+N-2}\cdots H_{k}|_{\infty}\leq\sqrt{d}|H_{k+N-1}H_{k+N-2}\cdots H_{k}|^{2}_{2}\leq\sqrt{d}\exp(-N\alpha h)\,.$
@@
for any $k,N\in\mathbb{N}$. When $2q^{N}\sqrt{s_{r_{N}}}<1$, we can let $k\to\infty$ and get the final bound in (2.2). The complete proof for Theorem 2.2 can be found in Appendix C. ∎
@@
(2.4) $\sqrt{\mathbb{E}_{\pi}[|\nabla V|_{\infty}^{2}]}\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\sqrt{\beta\log(2d)}\,.}$
@@
Under the assumption, it holds that $\nabla V$ is $\sqrt{\beta}$ sub-Gaussian under $\pi$ [1, Theorem 1.2]. The expected $\ell^{\infty}$ norm of a sub-Gaussian vector is bounded using standard results. The complete proof can be found in Appendix C.3. ∎
@@
Let Assumptions 1.1 and 2.1 hold with $s_{k}\leq C(k+1)^{n}$ for any $k\in\mathbb{N}$, where $C>0$ is a generic constant and $n\in\mathbb{N}$. Then for $h\leq 1/\beta$, we have
@@
The current upper bound in Theorem 2.4 involves $\sqrt{h}$. This is in alignment with the $W_{2}$ bound proved in the literature (under Assumption 1.1) based on coupling arguments where such square root dependence manifests. The second part of the bound, $\sqrt{h\log(2d)}O\left(\frac{\beta}{\alpha}\sqrt{d}\right)$, is consistent, up to logarithmic terms in $d$, with the established result for the $W_{2}$ bound. In fact, using the $W_{2}$ bound, we have that $W_{2,\ell^{\infty}}(\pi_{h},\pi)\leq W_{2}(\pi_{h},\pi)=O(\frac{\beta}{\alpha}\sqrt{dh})$.
[…]